OpenStudy (anonymous):

solve using the addition method 6x + 3y = 0 8x + 5y = 8

OpenStudy (anonymous):

5(6x+3y)=5*0 30x+15y=0 -3(8x+5y)=-3*8 -24x-15y=-24 30x+15y-24x-15y=0-24 6x=-24 x=-4 6*(-4)+3y=0 -24+3y=0 3y=24 y=8